OpenStudy (amenah8):

OPTIMIZATION: A wire 2 meters long is cut into two pieces. One piece is bent into a square for a frame for a stained glass ornament, while the other piece is bent into a circle for a TV antenna. To reduce storage space, where should the wire be cut to minimize the total area of both figures? Where should the wire be cut to maximize the total area?

OpenStudy (amenah8):

the length of wire used for: square = 8/(pi+4) m circle = 2-8/(pi+4) m

OpenStudy (amenah8):

but now i want the MAXIMUM for each

OpenStudy (bonnieisflash1.0):

@osprey is out of opensudy

OpenStudy (bonnieisflash1.0):

he left

OpenStudy (bonnieisflash1.0):

i know ho can help though @mathmate

OpenStudy (bonnieisflash1.0):

can you help @mathmate

OpenStudy (mathmate):

@Amenah8 are you there?

OpenStudy (bonnieisflash1.0):

i think so

OpenStudy (mathmate):

@bonnieisflash1.0 is your alias Amenah8 ?

OpenStudy (bonnieisflash1.0):

@mathmate is it all right to call you puppet master

OpenStudy (bonnieisflash1.0):

and NO

OpenStudy (mathmate):

@Amenah8 You can tag me when you're online. Good night!

OpenStudy (bonnieisflash1.0):

she's here

OpenStudy (amenah8):

@mathmate sorry my computer messed up!

satellite73 (satellite73):

maximum area?

OpenStudy (bonnieisflash1.0):

my computer is slow

OpenStudy (mathmate):

@Amenah8 no problem. You want the maximum total area, right?

OpenStudy (bonnieisflash1.0):

bye

satellite73 (satellite73):

i have a guess

satellite73 (satellite73):

@mathmate you got a guess?

OpenStudy (mathmate):

Nope, tell me about it! @satellite73

OpenStudy (amenah8):

the question was Where should the wire be cut to maximize the total area? so i guess the max for each, the square and the circle

satellite73 (satellite73):

my guess is to maximize, use all for the circle should get the biggest area that way

satellite73 (satellite73):

but if you did the "min" part you already have the function and the derivative and the min, use the other critical point if there is one, and also try the endpoints

OpenStudy (amenah8):

what do u mean?

satellite73 (satellite73):

you want to maximize and area with a fixed perimeter, make a circle it gives the biggest area

OpenStudy (mathmate):

@satellite73 Agree! If we split it up more or less half/half, it would be min.

satellite73 (satellite73):

did you get a function for the combined areas?

OpenStudy (mathmate):

@satellite73 I do, not sure if @Amenah8 did.

OpenStudy (mathmate):

@Amenah8 Have you been following @satellite73 posts?

OpenStudy (mathmate):

BTW, @Amenah8 Have you done calculus yet?

satellite73 (satellite73):

if you got some solution you must have a a function that gave the area if some amount was used for the circle] otherwise you couldn't have done it

OpenStudy (mathmate):

@Amenah8 Are you there at the keyboard? @satellite73 Yeah, the square takes slightly more wire than the circle.

OpenStudy (mathmate):

@satellite73 But perhaps OP has not done calculus. The question could be just to illustrate that Area is proportional to linear dimensions squared.

OpenStudy (amenah8):

no i'm in calculus now. it's supposed to be optimization --> i have the mins but i still need the max

OpenStudy (amenah8):

hold on i can paste my work when i found the min

OpenStudy (mathmate):

Good, read through the thread, you'll find the answer from @satellite73 's posts. If you have questions as to why, tag @satellite73 or me, your choice.

OpenStudy (amenah8):

derivative was (1/8)x - (1/2pi)(2-x)

OpenStudy (amenah8):

and square min was 8/(pi+4) and circle min was (2-8)/(pi+4)

OpenStudy (amenah8):

but how do i use that to find the max for each?

OpenStudy (amenah8):

@satellite73

OpenStudy (amenah8):

you said to split them in half?

satellite73 (satellite73):

no check the endpoints of the interval of your function, which i still don't see

satellite73 (satellite73):

i see "derivative was (1/8)x - (1/2pi)(2-x)" but i don;t see what the original function was

OpenStudy (amenah8):

oh hold on

satellite73 (satellite73):

oh here we don't need it

OpenStudy (amenah8):

A for circle and square combined = (1/16)x^2 + pi(2-x / 2pi)^2

satellite73 (satellite73):

the critical points are a) the zero of the derivative of b) where the derivative is undefined, i.e where the denominator is 0`

satellite73 (satellite73):

or the endpoints of the interval of the domain namely \(0\) and \(2\)

OpenStudy (amenah8):

so plug those into the original equation?

OpenStudy (mathmate):

@satellite73 For your info, both expressions for area and derivatives correspond to mine.

satellite73 (satellite73):

ok so good enough , check the endpoints of the interval, which are also where the derivative is undefined my guess is the one that says "all circle" is the biggest

OpenStudy (mathmate):

@satellite73 Right! http://prnt.sc/czfn0f

OpenStudy (amenah8):

ohhhh this makes much more sense!

OpenStudy (amenah8):

but it does not show a max

OpenStudy (mathmate):

Remember \(global\) max can occur at any of the critical points, or the closed endpoints of the domain.

OpenStudy (amenah8):

so about when y = 2.5?

satellite73 (satellite73):

is this your original function (1/16)x^2 + pi(2-x / 2pi)^2

OpenStudy (amenah8):

yes

satellite73 (satellite73):

is it clear what the endpoints of the interval are?

satellite73 (satellite73):

i mean endpoints of the domain

OpenStudy (skullpatrol):

Welcome back @satellite73 :D

satellite73 (satellite73):

thanks !

OpenStudy (amenah8):

wasn't the endpoint 0 and 2?

satellite73 (satellite73):

yes

satellite73 (satellite73):

check them

OpenStudy (amenah8):

.32 at 0 and .25 at 2

satellite73 (satellite73):

then 3.2 it is

OpenStudy (amenah8):

and that is the max for the square or the circle?

satellite73 (satellite73):

seems wrong to me though if you make a circle of radius two the area is \(4\pi\)

satellite73 (satellite73):

it is not the max for a square and a circle it is the max combined area

OpenStudy (amenah8):

so then what is the max for each separately?

OpenStudy (amenah8):

@satellite73 sorry my computer messed up

OpenStudy (mathmate):

@Amenah8 `.32 at 0 and .25 at 2` are the 'total' areas. The first one refers to x=0 (left extreme) where the size of the square is zero. So the circle gets a \(circumference\) of 2 metres. From there, you would extract the radius, given by 2\(\pi\)r=2, or r=1/\(\pi\) to give an area of \(A_{total}=0+\pi(1/\pi)^2\)=0.32. Similarly, 0.25 is at x=2, giving a circle of 0 circumference and a square of 2 metres in perimeter. The side is therefore 2/4=0.5, hence \(A_{total}=(1/2)^2+0=0.25\) It's up to you to choose the global maximum from all of the above, possibly with the help of the graph. Mathematically it should be the maximum among all critical points and end-points of domains (if applicable). Hope that helps to clear up queries.